OpenStudy (anonymous):

If u=1 and v=5, and the vectors make the angles of 130 and 180 degrees with the x-axis respectively, find the component form of the sum of u and v.

OpenStudy (dumbcow):

First find the x,y components for vectors u and v, then you can just add the vectors together by adding their respective components to go from polar to coordinate: x = r*cos(theta) y= r*sin(theta) where r is the magnitude of each vector, 1 and 5 respectively

OpenStudy (earthcitizen):

use de moivre's theorem

OpenStudy (anonymous):

de moivre's theorem should do the trick

OpenStudy (earthcitizen):

u=1+j0, v=5+j0

OpenStudy (dumbcow):

how does de'Moivres theorem apply to adding vectors

OpenStudy (earthcitizen):

\[like this z=x+jy =r(\cos(\theta)+jsin(\theta))\]

OpenStudy (earthcitizen):

\[u=1+j0 = r=\sqrt{1^2+0^2}= 1\]

OpenStudy (earthcitizen):

\[\theta=130 \]

OpenStudy (earthcitizen):

\[\theta=\tan^{-1} (y/x) = \tan^{-1} (0/1)+\pi\]

OpenStudy (earthcitizen):

r<0 = 1<180

OpenStudy (earthcitizen):

\[v= 5+j0 = r=\sqrt{25}=5\]

OpenStudy (earthcitizen):

\[\theta= 180\]

OpenStudy (earthcitizen):

yo ?

OpenStudy (earthcitizen):

\[u=1(\cos(130)+jsin(130)) =-0.643+j0.77\]

OpenStudy (earthcitizen):

\[v=5(\cos(180)+jsin(180))= -5+0=-5\]

OpenStudy (earthcitizen):

u+v =-5.643+j0.77

OpenStudy (earthcitizen):

are u sure the angles are for u and v respectively ?

OpenStudy (dumbcow):

ok basically same thing as converting the coordinates x = r*cos(theta) y= r*sin(theta) u = <cos(130), sin(130)> v = <-5, 0> u+v = <-5+cos(130), sin(130)> = -5.63i +0.77j

OpenStudy (earthcitizen):

is i not the same as j ?

OpenStudy (earthcitizen):

u=1<130, v=5<180 in polar form

OpenStudy (dumbcow):

i is x-part, j is the y-part, when talking about rectangular coordinates